# Cuipo Jiang and Jingtian Song

### v2

@@ 1 Introduction
for $a,b\in V$, where $\bar{a}:=a+C_{2}(V)$ is the image of $a$ in $R_{V}$. The associated variety $X_{V}$ of $V$ is the reduced scheme $X_{V}=\operatorname{Specm}(R_{V})$ of $R_{V}$ [Ara12]. This fundamental invariant of $V$ has been extensively studied in [BFM91, Zhu96, ABD04, Miy04, Ara12, Ara15a, Ara15b, AM18a, AK18, AM17, AJM21, ADF+24, AFK24].
@@
the variety $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit in $\mathfrak{g}$. In particular, if $k$ is a non-degenerate admissible number, that is,
@@
then $X_{L_{k}(\mathfrak{g})}$ is the nilpotent cone ${\mathcal{N}}(\mathfrak{g})$ [Ara15a]. If $k=-h^{\vee}$, it follows from [FF92, FG04] that $X_{L_{k}(\mathfrak{g})}$ is also the nilpotent cone. It was observed in [AM18a, AM18b, AM17, ADF+24] that there are cases when $L_{k}(\mathfrak{g})$ is non-admissible and $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit, which provide examples for the conjecture in physics [XY21] that, in view of the 4D/2D duality, there should be a large list of non-admissible simple affine vertex algebras whose associated varieties are the closures of some nilpotent orbits. There are also cases where $X_{L_{k}(\mathfrak{g})}$ is neither $\mathfrak{g}^{*}$ nor contained in the nilpotent cone ${\mathcal{N}}(\mathfrak{g})$. It was proved in [AM17] that $X_{L_{-1}(sl_{n})}$, $n\geqslant 4$, $X_{L_{-m}(sl_{2m})}$, $m\geqslant 2$, $X_{L_{2-r}(so_{2r})}$, $r\in 2{\mathbb{Z}}_{+}+1$, are closures of sheets. In [AFK24], it was shown that the associated variety of a simple affine vertex algebras is contained in the closure of the Dixmier sheet when a chiralization of generalized Grothendieck’s simultaneous resolution exists. The variety $X_{L_{-\frac{5}{2}}}(sl_{4})$ of $L_{-\frac{5}{2}}(sl_{4})$ was given in [Fas22], which turns out to be the closure of a Jordan class of $sl_{4}$.
In general, the problem of determining $X_{L_{k}(\mathfrak{g})}$ is widely open. Up to now, the variety $X_{L_{k}(\mathfrak{g})}$ has been completely determined for all $k\in{\mathbb{C}}$ only when $\mathfrak{g}=sl_{2}$. In this paper, we will determine $X_{L_{k}(sl_{3})}$ for all non-admissible numbers $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$. Then together with the known results $X_{L_{k}(\mathfrak{g})}$ for the critical case and admissible cases given in [FF92, FG04, Ara12], we will obtain a complete characterization of $X_{L_{k}(sl_{3})}$ for all $k\in{\mathbb{C}}$.
Another question we are concerned about in this paper is the characterization of the maximal ideal of $V^{k}(sl_{n})$, where $k=-n+\frac{n-1}{q}$, $n\geqslant 3$, $q\geqslant 2$, $(n-1,q)=1$. If $q=1$, that is, $k=-1$, the maximal ideal of $V^{-1}(sl_{3})$ is generated by two singular vectors of conformal weight 3 [AP08, AKMF+20], and the maximal ideal of $V^{-1}(sl_{n})$ is generated by one singular vector of conformal weight 2 [AP08, AM17]. By using the character formulas given in [KT00], the fact that the coefficients as values of Kazhan-Lusztig (or inverse KL) polynomials at 1 in the formulas depend only on the corresponding Coxeter group ( see also Theorem 11 in [Fie06] from the category point of view), and the results for the $k=-1$ case in [KW18, BKK24, AP08, AM17, AKMF+20], we prove our first main result as follows (also see Theorem 3.8).
@@
where ${\mathfrak{l}}_{1}=\mathfrak{h}\oplus{\mathbb{C}}e_{\theta}\oplus{\mathbb{C}}f_{\theta}$, $\lambda=h_{1}-h_{2}$, and ${\mathcal{O}}_{\min}$ is the minimal non-zero nilpotent orbit of $sl_{3}({\mathbb{C}})$. In particular, $X_{L_{-1}(sl_{3})}$ is irreducible and $\dim X_{L_{-1}(sl_{3})}=5$.
@@
Representations of $L_{k}(sl_{3})$ have been extensively studied in [AP08, Ada16, AFR17, Cre18, FLLZ18, AM21, AK21, AKR21, FKR21, FR22, KRW22, AFP23a, CRR23, ACG24], etc. We here have the following conjecture for $L_{k}(sl_{3})$ (also see Conjecture 4.11).
@@
Let $f$ be a minimal nilpotent element of $\mathfrak{g}=sl_{3}$. If $k=-3+\frac{2m+1}{2}$, $m\geqslant 1$, it was shown in [Ara13] that $X_{W_{k}(sl_{3},f)}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational. If $k=-1$, it has been proved in [AKF+18] that $W_{-1}(sl_{3},f)$ is isomorphic to the rank one Heisenberg vertex operator algebra. So $X_{W_{-1}(sl_{3},f)}$ is one-dimensional. We have the following result for $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$ (also see Theorem 5.3).
@@
Let $f$ be a regular nilpotent element of $\mathfrak{g}=sl_{3}$. If $k=-3+\frac{p}{q}$, $p,q\geqslant 3$, $(p,q)=1$, by [Ara15b], $X_{W_{k}(sl_{3})}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational. Our fifth main result is given as follows (also see Theorem 5.4).
@@
In particular, $\dim X_{W_{k}(sl_{3})}=1$. Furthermore, $W_{k}(sl_{3},f)$ is not quasi-lisse.
The rest of this paper is organized as follows. In Section 2, we recollect some concepts and results that will be needed later. In Section 3, we study singular vectors and maximal ideals of $V^{k}(sl_{n})$ for $k=-n+\frac{n-1}{q}$ with $n\geqslant 3$, $q\geqslant 1$, $(n-1,q)=1$. Section 4 is dedicated to determine the associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$. In Section 5, we obtain the associated varieties of the simple affine $W$-algebras $W_{k}(sl_{3},f)$.
@@ Acknowledgements
We appreciate Tomoyuki Arakawa, Dražen Adamović, and Anne Moreau for valuable discussions and comments. In particular, the first author is grateful to Dražen Adamović for showing her the very helpful result of Theorem 8.2 in his paper [AFP23b] joint with Pierluigi Möseneder Frajria and Paolo Papi, and to Tomoyuki Arakawara for telling her the papers [Fie06] and [BKK24], and for the discussions.
@@ 2.1 $C_{2}$-algebras and associated varieties
For a vertex algebra $V$, define $C_{2}(V)=\operatorname{span}_{{\mathbb{C}}}\{a_{-2}b\mid a,b\in V\}$ and
@@
Let $V$ be a vertex algebra. Then $V$ is called quasi-lisse if $X_{V}$ has finitely many symplectic leaves.
@@
Let $V$ be a finitely strongly generated vertex algebra. If $V$ is quasi-lisse, then $V$ has finitely many ordinary modules.
@@ 2.2 Affine vertex algebras
The space $V^{k}(\mathfrak{g})$ is naturally graded,
@@ 2.3 Zhu’s algebra of affine vertex algebras
For $\lambda\in\mathfrak{h}^{*}$, let $L(\lambda)$ be the irreducible module of $V^{k}(\mathfrak{g})$ induced from the simple highest weight $\mathfrak{g}$-module $L(\bar{\lambda})$. Then by [Zhu96], $L(\lambda)$ is a module of $L_{k}(\mathfrak{g})$ if and only if $I_{k}(L(\bar{\lambda}))=0$. We have the following result from [Ara16].
@@ 2.5 Zhu’s $C_{2}$-algebras and associated varieties of affine vertex algebras
2.6 Sheet and Jordan classes
@@ 2.7 Affine $W$-algebras
Here $H^{\bullet}_{DS,f}(M)$ denotes the BRST cohomology of the generalized quantized Drinfeld-Sokolov reduction associated with $f\in\mathcal{N}(\mathfrak{g})$ with coefficients in a $V^{k}(\mathfrak{g})$-module $M$ [KRW03]. Recall that we have a natural isomorphism $R_{W^{k}(\mathfrak{g},f)}\cong{\mathbb{C}}[\mathscr{S}_{f}]$ of Poisson algebras [DSK06, Ara15a], so that
@@
    $H_{DS,f_{\theta}}^{i}(M)=0$ for all $i\neq 0$, $M\in{\mathcal{O}}_{k}$. In particular, the functor
@@
    $H_{DS,f}^{i}(M)=0$ for all $i\neq 0$, $M\in\operatorname{KL}_{k}$. In particular, the functor
@@
$r(k+h^{\vee})\in{\mathbb{Q}}_{\geqslant 0}\backslash\left\{\frac{1}{m}\mid m\in{\mathbb{Z}}_{\geqslant 1}\right\},$ (3.1)
where $r$ is the lacing number of $\mathfrak{g}$.
@@
In this section, we will prove that, for $k=-3+\frac{2}{2m+1}$, $m\in{\mathbb{Z}}_{\geqslant 1}$, the maximal ideal of $V^{k}(sl_{3})$ is generated by two singular vectors of conformal weight $3(2m+1)$, and for $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$, the maximal ideal of $V^{k}(sl_{n})$ is generated by a singular vector of conformal weight $2q$. We reach our results by using the Kashiwara-Tanisaki characters [KT00] (see also [Fie06]), the character formulas for the $k=-1$ case given in [KW18, BKK24], and the results established in [AP08, AKMF+20, AM17].
@@ 3.1 Kashiwara-Tanisaki Characters
For $w\in\widehat{W}(\lambda)$, denote by $l_{\lambda}(w)$ the length of $w$. Denote the Bruhat ordering of $\widehat{W}(\lambda)$ by $\geqslant_{\lambda}$. For $y,w\in\widehat{W}(\lambda)$, denote by $P^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the corresponding Kazhdan-Lusztig polynomial [KL79], and by $Q^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the inverse Kazhdan-Lusztig polynomial defined by
@@
For $\lambda\in\widehat{\mathfrak{h}}^{*}$, let $M(\lambda)$ (resp. $L(\lambda)$) be the Verma module (resp. simple module) of $\widehat{\mathfrak{g}}$ with highest weight $\lambda$. We have the following result from [KT00].
@@
$\sigma s_{\alpha_{0}}=s_{\beta_{0}},\sigma(s_{\alpha_{i}})=s_{\alpha_{i}},\ i=1,2,\cdots,n-1.$
@@
for $\lambda\in\widehat{\mathfrak{h}}^{*}$. Recall from [Kac90], for any $x\in\widehat{W}$, there exists $y\in W$ and $\gamma\in Q$ such that $x=yt_{\gamma}$. We have the following lemma.
@@
Let $n\geqslant 3$, and let $L(\Lambda)$ be an irreducible level $k=-1$ $\widehat{sl}_{n}$-module with highest weight
@@
$\widehat{R}{\rm ch}L(\Lambda)=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1}(resp.\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}({\Lambda+\widehat{\rho}})-\widehat{\rho}}.$
@@
$\widehat{R}{\rm ch}L(-\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
We also need the following results from [KK79], [DGK82], [Kum87], [KW89].
Lemma 3.6.
Let $\lambda\in{\mathcal{C}}^{+}$, and $w\in\widehat{W}(\lambda)$. If $L(\mu)$ is a subquotient of $M(w\circ\lambda)$, then there exists $y\in\widehat{W}(\lambda)$ such that $\mu=yw\circ\lambda$.
Recall the following results from [AP08], [AKMF+20], and [AM17].
@@
    $u=e_{\theta}(-1)e_{\alpha_{2}+\cdots\alpha_{n-2}}(-1){\bf 1}-e_{\alpha_{1}+\cdots\alpha_{n-2}}(-1)e_{\alpha_{2}+\cdots\alpha_{n-1}}(-1){\bf 1}.$
@@
    [AKMF+20] The ideal of $V^{-1}(sl_{3})$ generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$.
@@
Let $\mathfrak{g}=sl_{n}$, $n\geqslant 3$. Notice that if $q=1$, then $k=-1$ and $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))=(\widehat{W},S)$. So $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))\cong(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2,(q,n-1)=1$. By Lemma 3.3, and the definitions of $P^{\lambda}_{y,w}(q)$ and $Q^{\lambda}_{y,w}(q)$ [KL79], we have for $y,w\in\widehat{W}$,
@@
$l_{-\Lambda_{0}}(y)=l_{k\Lambda_{0}}\sigma(y),\ l_{-\Lambda_{0}}(w)=l_{k\Lambda_{0}}\sigma(w),$ (3.10)

### v3

# Associated varieties of simple affine VOAs $L_{k}(sl_{3})$ and $W$-algebras $W_{k}(sl_{3},f)$

@@ 1 Introduction
for $a,b\in V$, where $\bar{a}:=a+C_{2}(V)$ is the image of $a$ in $R_{V}$. The associated variety $X_{V}$ of $V$ is the reduced scheme $X_{V}=\operatorname{Specm}(R_{V})$ of $R_{V}$ [Ara12]. This fundamental invariant of $V$ has been extensively studied in [BFM91, Zhu96, ABD04, Miy04, Ara12, Ara15a, Ara15b, AM18a, AK18, AM17, AJM21, ADF+25, AFK24].
@@
the variety $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit in $\mathfrak{g}$ [Ara15a]. In particular, if $k$ is a non-degenerate admissible number, that is,
@@
then $X_{L_{k}(\mathfrak{g})}$ is the nilpotent cone ${\mathcal{N}}(\mathfrak{g})$ [Ara15a]. If $k=-h^{\vee}$, it follows from [FF92, FG04] that $X_{L_{k}(\mathfrak{g})}$ is also the nilpotent cone. It was observed in [AM18a, AM18b, AM17, ADF+25] that there are cases when $L_{k}(\mathfrak{g})$ is non-admissible and $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit, which provide examples for the conjecture in physics [XY21] that, in view of the 4D/2D duality, there should be a large list of non-admissible simple affine vertex algebras whose associated varieties are the closures of some nilpotent orbits. There are also cases where $X_{L_{k}(\mathfrak{g})}$ is neither $\mathfrak{g}^{*}$ nor contained in the nilpotent cone ${\mathcal{N}}(\mathfrak{g})$. It was proved in [AM17] that $X_{L_{-1}(sl_{n})}$, $n\geqslant 4$, $X_{L_{-m}(sl_{2m})}$, $m\geqslant 2$, $X_{L_{2-r}(so_{2r})}$, $r\in 2{\mathbb{Z}}_{+}+1$, are closures of sheets. In [AFK24], it was shown that the associated variety of a simple affine vertex algebras is contained in the closure of a Dixmier sheet when a chiralization of generalized Grothendieck’s simultaneous resolution exists. The variety $X_{L_{-\frac{5}{2}}(sl_{4})}$ of $L_{-\frac{5}{2}}(sl_{4})$ was given in [Fas22], which turns out to be the closure of a Jordan class of $sl_{4}$.
In general, the problem of determining $X_{L_{k}(\mathfrak{g})}$ is widely open. Up to now, the variety $X_{L_{k}(\mathfrak{g})}$ has been completely determined for all $k\in{\mathbb{C}}$ only when $\mathfrak{g}=sl_{2}$. In this paper, we will determine $X_{L_{k}(sl_{3})}$ for all non-admissible numbers $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$. Then together with the results $X_{L_{k}(\mathfrak{g})}$ for the critical case and admissible cases given in [FF92, FG04, Ara12], we will obtain a complete characterization of $X_{L_{k}(sl_{3})}$ for all $k\in{\mathbb{C}}$.
Another question we are concerned about in this paper is the characterization of the maximal ideal of $V^{k}(sl_{n})$, where $k=-n+\frac{n-1}{q}$, $n\geqslant 3$, $q\geqslant 2$, $(n-1,q)=1$. If $q=1$, that is, $k=-1$, the maximal ideal of $V^{-1}(sl_{3})$ is generated by two singular vectors of conformal weight 3 [AP08, AKMP+20], and the maximal ideal of $V^{-1}(sl_{n})$ for $n\geqslant 4$ is generated by one singular vector of conformal weight 2 [AP08, AM17]. By using the character formulas given in [KT00], the fact that the coefficients as values of Kazhan-Lusztig (or inverse KL) polynomials at 1 in the formulas depend only on the corresponding Coxeter group, the category equivalences given in [Fie06], and the results for the $k=-1$ case in [KW18, BKK24, AP08, AM17, AKMP+20], we prove our first main result as follows (also see Theorem 3.11).
@@
where ${\mathfrak{l}}_{1}=\mathfrak{h}\oplus{\mathbb{C}}e_{\theta}\oplus{\mathbb{C}}f_{\theta}$, $\lambda=h_{1}-h_{2}$, and ${\mathcal{O}}_{\min}$ is the minimal non-zero nilpotent orbit of $sl_{3}$. In particular, $X_{L_{-1}(sl_{3})}$ is irreducible and $\dim X_{L_{-1}(sl_{3})}=5$.
Remark 1.3.
The result of Theorem 1.2 is also covered in [AFK24] in a different and more conceptual way.
@@
Representations of $L_{k}(sl_{3})$ have been extensively studied in [AP08, Ada16, AFR17, Cre18, FLLZ18, AM21, AK21, AKR21, FKR21, FR22, KRW22, AMP23, CRR23, ACG24, AKR24], etc. We have the following conjecture for $L_{k}(sl_{3})$ (also see Conjecture 4.11).
@@
Let $f$ be a minimal nilpotent element of $\mathfrak{g}=sl_{3}$. If $k=-3+\frac{2m+1}{2}$, $m\geqslant 1$, it was shown in [Ara13] that $X_{W_{k}(sl_{3},f)}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational. If $k=-1$, it has been proved in [AKMP+18] that $W_{-1}(sl_{3},f)$ is isomorphic to the rank one Heisenberg vertex operator algebra. So $X_{W_{-1}(sl_{3},f)}$ is one-dimensional. We have the following result for $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$ (also see Theorem 5.3).
@@
Let $f$ be a regular nilpotent element of $\mathfrak{g}=sl_{3}$. If $k=-3+\frac{p}{q}$, $p,q\geqslant 3$, $(p,q)=1$, by [Ara15b], $X_{W_{k}(sl_{3},f)}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational. Our fifth main result is given as follows (also see Theorem 5.4).
@@
$X_{W_{k}(sl_{3},f)}=\left\{\left[\begin{array}[]{ccc}0&\frac{3}{4}\mu^{2}&\frac{1}{2}\mu^{3}\\ 2&0&\frac{3}{4}\mu^{2}\\ 0&2&0\end{array}\right]|~{}\mu\in{\mathbb{C}}\right\}.$
In particular, $\dim X_{W_{k}(sl_{3},f)}=1$. Furthermore, $W_{k}(sl_{3},f)$ is not quasi-lisse.
The rest of this paper is organized as follows. In Section 2, we recollect some concepts and results that will be needed later. In Section 3, we study singular vectors and maximal ideals of $V^{k}(sl_{n})$ for $k=-n+\frac{n-1}{q}$ with $n\geqslant 3$, $q\geqslant 1$, $(n-1,q)=1$. Section 4 is dedicated to determining the associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$. In Section 5, we obtain the associated varieties of the simple affine $W$-algebras $W_{k}(sl_{3},f)$.
@@ Acknowledgements
We appreciate Tomoyuki Arakawa, Dražen Adamović, and Anne Moreau for invaluable discussions and comments. In particular, the first author is grateful to Dražen Adamović for showing her the very helpful result of Theorem 8.2 in his paper [AMP23] joint with Pierluigi Möseneder Frajria and Paolo Papi, and to Tomoyuki Arakawara for telling her the papers [Fie06] and [BKK24], and for the discussions. Results in this paper were presented in part in conferences “Representation Theory XVIII” in Dubrovnik, 2023; “Vertex Algebras , Infinite Dimensional Lie Algebras, and Related Topics” in Rome, 2023; “RIMS Joint Research Activity: Vertex Algebras and Related Topics ” in Kyoto, 2024; and “TYMRC-Workshop on Quantum Algebras and Representations” in Shenzhen, 2024. We thank the organizers of these conferences. Finally, we are very grateful to the referees for invaluable comments and suggestions.
@@ 2.1 $C_{2}$-algebras and associated varieties
For a vertex algebra $V$, let $C_{2}(V)=\operatorname{span}_{{\mathbb{C}}}\{a_{-2}b\mid a,b\in V\}$ and
@@
Let $V$ be a vertex algebra. $V$ is called quasi-lisse if $X_{V}$ has finitely many symplectic leaves.
@@
Let $V$ be a finitely strongly generated vertex algebra. If $V$ is quasi-lisse, then $V$ has finitely many irreducible ordinary modules.
@@ 2.2 Affine vertex algebras
The space $V^{k}(\mathfrak{g})$ is naturally graded as follows:
@@ 2.3 Zhu’s algebra of affine vertex algebras
For $\lambda\in\mathfrak{h}^{*}$, let $L(\lambda)$ be the irreducible module of $V^{k}(\mathfrak{g})$ induced from the simple highest weight $\mathfrak{g}$-module $L(\bar{\lambda})$. Then by [Zhu96], $L(\lambda)$ is a module of $L_{k}(\mathfrak{g})$ if and only if $I_{k}$ acts on $L(\bar{\lambda})$ as zero. We have the following result from [Ara16].
@@ 2.5 Zhu’s $C_{2}$-algebras and associated varieties of affine vertex algebras
2.6 Sheets and Jordan classes
@@ 2.7 Affine $W$-algebras
Here $H^{\bullet}_{DS,f}(M)$ denotes the BRST cohomology of the generalized quantized Drinfeld-Sokolov reduction associated to $f\in\mathcal{N}(\mathfrak{g})$ with coefficients in a $V^{k}(\mathfrak{g})$-module $M$ [KRW03]. Recall that we have a natural isomorphism $R_{W^{k}(\mathfrak{g},f)}\cong{\mathbb{C}}[\mathscr{S}_{f}]$ of Poisson algebras [DSK06, Ara15a], so that
@@
    $H_{DS,f_{\theta}}^{i}(M)=0$ for all $i\neq 0$, and $M\in{\mathcal{O}}_{k}$. In particular, the functor
@@
    $H_{DS,f}^{i}(M)=0$ for all $i\neq 0$, and $M\in\operatorname{KL}_{k}$. In particular, the functor
@@
$r^{\vee}(k+h^{\vee})\in{\mathbb{Q}}_{\geqslant 0}\backslash\left\{\frac{1}{m}\mid m\in{\mathbb{Z}}_{\geqslant 1}\right\},$ (3.1)
where $r^{\vee}$ is the lacing number of $\mathfrak{g}$.
@@
In this section, we will prove that, for $k=-3+\frac{2}{2m+1}$, $m\in{\mathbb{Z}}_{\geqslant 1}$, the maximal ideal of $V^{k}(sl_{3})$ is generated by two singular vectors of conformal weight $3(2m+1)$, and for $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$, the maximal ideal of $V^{k}(sl_{n})$ is generated by a singular vector of conformal weight $2q$. We reach our results by using the Kashiwara-Tanisaki characters [KT00], category equivalences established in [Fie06], the character formulas for the $k=-1$ case given in [KW18, BKK24], and the results established in [AP08, AKMP+20, AM17].
@@ 3.1 Kashiwara-Tanisaki Characters
For $w\in\widehat{W}(\lambda)$, denote by $l_{\lambda}(w)$ the length of $w$. Denote the Bruhat ordering of $\widehat{W}(\lambda)$ by $\geqslant_{\lambda}$. For $y,w\in\widehat{W}(\lambda)$, denote by $P^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the associated Kazhdan-Lusztig polynomial [KL79], and by $Q^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the inverse Kazhdan-Lusztig polynomial defined by
@@
For $\lambda\in\widehat{\mathfrak{h}}^{*}$, let $M(\lambda)$ (resp. $L(\lambda)$) be the Verma module (resp. simple module) of $\widehat{\mathfrak{g}}$ with highest weight $\lambda$. The following lemma comes from [KK79], [DGK82], [Kum87], [KW89].
Lemma 3.2.
Let $\lambda\in{\mathcal{C}}^{+}$, and $w\in\widehat{W}(\lambda)$. If $L(\mu)$ is a subquotient of $M(w\circ\lambda)$, then there exists $y\in\widehat{W}(\lambda)$ such that $\mu=yw\circ\lambda$, and $w\leqslant_{\lambda}yw$.
We have the following character formulas from [KT00].
@@
We have the following lemma from [KT98], [Fie06].
Lemma 3.4.
Let $\lambda\in{\cal C}^{+}$ or ${\cal C}^{-}$, and $w,w^{\prime}\in\widehat{W}(\lambda)$ be such that $w^{\prime}\cdot\lambda\leqslant w\cdot\lambda$. Then
${\rm dim}{\rm Hom}_{\mathbb{C}}(M(w^{\prime}\cdot\lambda),M(w\cdot\lambda))=1.$ (3.2)
The following theorem comes from [Fie06].
Theorem 3.5.
Let ${\mathfrak{h}}\subset{\mathfrak{b}}\subset{\mathfrak{g}}$ and ${\mathfrak{h}}^{\prime}\subset{\mathfrak{b}}^{\prime}\subset{\mathfrak{g}}^{\prime}$ be symmetrizable Kac-Moody Lie algebras with Cartan and Borel subalgebras, respectively. Let $\Lambda\in\mathfrak{h}^{*}/\sim$ and $\Lambda^{\prime}\in(\mathfrak{h}^{\prime})^{*}/\sim^{\prime}$ be two equivalent classes outside the critical hyperplanes and let ${\cal O}_{\Lambda}$ and ${\cal O}^{\prime}_{\Lambda^{\prime}}$ be the corresponding indecomposable blocks, where $\sim$ and $\sim^{\prime}$ are the usual relation in $\mathfrak{h}^{*}$ and $(\mathfrak{h}^{\prime})^{*}$, respectively (see [KK79]). Suppose the following conditions hold:
  1. (1)
    There exist $\lambda\in\Lambda$ and $\lambda^{\prime}\in\Lambda^{\prime}$ which are either both dominant or antidominant;
  2. (2)
    $(W(\lambda),S(\lambda))\cong(W(\lambda^{\prime}),S(\lambda^{\prime}))$ are isomorphic as Coxeter groups;
  3. (3)
    $W_{0}(\lambda)\cong W_{0}(\lambda^{\prime})$ under the same isomorphism and both sets are finite.
Then there exists an equivalence of categories
${\cal O}_{\Lambda}\cong{\cal O^{\prime}}_{\Lambda^{\prime}}.$
Remark 3.6.
As an application of Theorem 3.5, the author in [Fie06] proved Theorem 3.3 for symmetrizable Kac-Moody Lie algbera $\mathfrak{g}$, and regular $\Lambda\in{\mathfrak{h}}^{*}/\sim$ satisfying the condition in Theroem 3.5 and that $(W(\Lambda),S(\Lambda))$ is finite or affine type, and dominant or antidominant $\lambda\in\Lambda$.
@@
$\sigma s_{\alpha_{0}}=s_{\beta_{0}},\sigma s_{\alpha_{i}}=s_{\alpha_{i}},\ i=1,2,\cdots,n-1.$
In particular, $\sigma(W_{0}(-\Lambda_{0}))=W_{0}(k\Lambda_{0})$.
@@
for $\lambda\in\widehat{\mathfrak{h}}^{*}$.
Recall from [Kac90], for any $x\in\widehat{W}$, there exists $y\in W$ and $\gamma\in Q$ such that $x=yt_{\gamma}$. We have the following lemma.
@@
Let $n\geqslant 3$, and let $L(\Lambda)$ be an irreducible $\widehat{sl}_{n}$-module at level $k=-1$ with highest weight
@@
$\widehat{R}{\rm ch}L(\Lambda)=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1}(resp.\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}({\Lambda+\widehat{\rho}})-\widehat{\rho}},$
@@
$\widehat{R}{\rm ch}L(-\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$ (3.9)
Recall the following results from [AP08], [AKMP+20], and [AM17].
@@
    $u=e_{\theta}(-1)e_{\alpha_{2}+\cdots+\alpha_{n-2}}(-1){\bf 1}-e_{\alpha_{1}+\cdots+\alpha_{n-2}}(-1)e_{\alpha_{2}+\cdots+\alpha_{n-1}}(-1){\bf 1}.$
@@
    [AKMP+20] The ideal of $V^{-1}(sl_{3})$ generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$.
@@
Let $\mathfrak{g}=sl_{n}$, $n\geqslant 3$. It is obvious that $-\Lambda_{0},k\Lambda_{0}\in{\mathcal{C}}^{+}$, for $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$.
Notice that if $q=1$, then $k=-1$ and $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))=(\widehat{W},S)$. So by Lemma 3.7, $(\widehat{W},S)\cong(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$ through $\sigma$, for $k=-n+\frac{n-1}{q}$, $q\geqslant 2,(q,n-1)=1$. In particular,
$\sigma(s_{\alpha_{0}})=s_{\beta_{0}},\ \sigma(s_{\alpha_{i}})=s_{\alpha_{i}},\ i=1,2,\cdots,n-1.$ (3.11)
Then for $w,y\in\widehat{W}$,